OpenStudy (anonymous):

Will someone check my answers? Multiply or Divide. 1. 5x+10/x+2 times 2x/4x-10 My answer for that one was (5x)/(2x-5) 2. 3n^2-n/n^2-1 divided by n^2/n+1 My answer was n(3n-1)/(n-1)

OpenStudy (anonymous):

re My answer for that one was (5x)/(2x-5) EXCELLENT

OpenStudy (anonymous):

for the second one I get this (3n^2 - 1) / (n(n-1)) how did u get urs?

OpenStudy (anonymous):

Honestly I have no idea, Algebra is my worst subject, and it confuses me so much. Thank you for checking for me.

OpenStudy (anonymous):

sorry I have a typo above

OpenStudy (anonymous):

its (3n-1) / (n(n-1))

OpenStudy (anonymous):

Okay thank you. :)

OpenStudy (anonymous):

u got the first right so u r good, not bad

OpenStudy (campbell_st):

your answer for the 1st problem is correct

OpenStudy (anonymous):

you only got the n in the wrong spot for the second answer, can you repeat ur work and see why?

OpenStudy (anonymous):

u have in the numerator and I have it in denominr

OpenStudy (campbell_st):

the 2nd question has an error. you are correct in factoring when you divide by a fraction, flip and multiply \[\frac{n(3n - 1)}{(n -1)(n + 1)} \times \frac{(n + 1)}{n^2}\] now remove the common factors leaving and see what you have left